SOHO DAILY REPORT Operations Day - 353/2001 (19 December 01) Mission Day 2210 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-011219 ===================================================================== PASS #1 Madrid DSS66 0619-1157 PASS #2 Goldstone DSS27 1329-2013 PASS #3 Canberra DSS46 2009-354/0020 ********************************************************************* PASS #1 Madrid DSS66 0619-1157 Anomalies: None NRT: None Activities Summary: - SSR Dump - Collected MDI-M (0731 - 0747) - Collected MDI-H (0741 - 1130) - Min/Max Data Collection (PROS) - OBT Distribution to SWAN, CELIAS - Turn OFF BCR1 (OCD#1597) - SSR SEF Reset - SSR To Record ------------------ PASS #2 Goldstone DSS27 1329-2013 Anomalies: None NRT: 1634 - 1652 Activities Summary: - SSR Dump - Collected MDI-M (1431 - 1632) - Collected MDI-H (1632 - 2013) - Uplinked delayed loads VIRA11220, SWA011218161 - Reaction Wheel Speed Monitoring Daily Adjustment - CRS Drift Estimation Update - SSR SEF Reset ------------------ PASS #3 Canberra DSS46 2009-354/0020 Anomalies: G01-0503 2021 12 minutes degraded TM. Antenna would not stay in autotrack mode. Data is nonrecoverable. DR#C100718. G01-0504 2114 Wayward command. 14 min commanding delay. 8 commands retransmitted. 7 TC Block sequence errors. G01-0505 2144 Wayward command. 20 min commanding delay. 19 commands retransmitted. 18 TC Block sequence errors. G01-0506 2158 Commands failed to reach station. SCP failure. 13 min commanding delay. DR#C100720. NRT: 2014 - 2114 2132 - 2144 2216 - 2231 Activities Summary: - Collected MDI-H (2009 - 0005) - Clear anomaly table - Turn BCR1 ON (OCD#1597) - SSR SEF Reset - SSR Record ------------------ -- End of Report --